please help..!
Find the slope-intercept form for the equation of the line which passes through the point (–1, 3) and has a slope of –2.
---
Find the slope-intercept form for the equation of the line which passes through the point (–3, 2) and the origin

Found 2 solutions by rfer, stanbon:
Answer by rfer(12662)

y=mx+b
3=-2*-1+b
3=2+b
b=1
--------
y=-2x+1

Answer by stanbon(57361)

Find the slope-intercept form for the equation of the line which passes through the point (–1, 3) and has a slope of –2.
---
Form: y = mx+b
You have y = 3 when x = -1 and you have m = -2
Solve for "b";
3 = -2*3+b
3 = -6+b
b = 9
---
Equation:
y = -2x+9
=================
---
Find the slope-intercept form for the equation of the line
which passes through the point (–3, 2) and the origin
---
You have y = 0 when x = 0
You have slope = (2-0)/(-3-0) = -2/3
-----
Solve for "b":
y = mx + b
0 = (-2/3)*0 + b
b = 0
----
Y = (-2/3)x
==================
Cheers,
Stan H.
====================